OpenStudy (anonymous):

Tickets to the zoo cost $14 for adults and $10 for children. A group of 12 people went to the zoo, and the tickets cost $140. The system of equations models this situation, where x is the number of adults and y is the number of children. x+y=12 14x+10y=140 How many adults and how many children were in the group? A. 5 adults and 7 children B. 6 adults and 6 children C. 7 adults and 5 children D. 8 adults and 4 children PLEASE HELP ASAP

OpenStudy (anonymous):

my mom said it's either B or C but i think it can also be D so not really sure

OpenStudy (anonymous):

Solve the following for a and c. \[\{14 a+10 c=140,a+c=12\} \] {a = 5, c = 7}

OpenStudy (anonymous):

\[x+y = 12\] \[14x+10y = 140\] You make \[[x + y = 12] *10 \] that is \[10x +10y = 120\] then you do 14x +10y = 140 - (10x +10y = 120) then you get 4x = 20 x = 5

OpenStudy (anonymous):

then x + y = 12 5 + y = 12 y = 7

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

You can also make x+y = 12 ----> x = 12 - y 14x + 10y = 144 --> using x = 12 -y then you get 14(12-y) + 10y = 140 168 -14y +10y = 140 168 - 140 = 14y -10y 28 = 4y y = 7 Answer is A

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

A.